clink12:

Write the equation of a line that is parallel to y=-2x + 4 and passes through the point (4, 0).

dude:

Parallel lines have the same slope, what is the slope of this line?

clink12:

I don't know.:|

dude:

The equation of a line in slop-intercept form is written as \(y=mx+b\) m is the slope and b is he y-intercept Do you now know what the slope is?

clink12:

-2x

dude:

Its just m but you're right its -2 So we make a new equation this way \(y=-2x+b\) We were given a point (4,0) so we substitute that point into the equation, do you know hot to do that?

clink12:

Substitute it for b?

dude:

Points are written as (x,y) :P

dude:

Substitute x and y

clink12:

So 0=-2*4 + 4?

dude:

b is not solved for \(0=-2*4+b\) You're solving for b